OpenStudy (anonymous):

How do I prove for any natural number (n), this expression is divisible by 10? (3^(n+2)-2^(n+2)+3^n-2^n)/10

OpenStudy (anonymous):

If you're proving for any natural number, then you're probably going to end up using a proof by induction. The base case is trivial, I haven't attempted the induction step yet.

OpenStudy (zarkon):

\[3^{n+2}-2^{n+2}+3^n-2^n\] \[=3^{n}3^2-2^{n}2^2+3^n-2^n\] \[=3^{n}9-2^{n}4+3^n-2^n\] \[=10\cdot3^{n}-5\cdot 2^{n}\] \[=10\cdot3^{n}-10\cdot 2^{n-1}\] \[=10(3^{n}- 2^{n-1})\]

OpenStudy (anonymous):

Huh. Easier than I thought.